The perimeter of a triangle is 57.The first side is twice the length of the second side.The third side is seven more than the second side.What is the length of each side?
Answer by Mathtut(3670)

A+B+C=57 A=2B and C=B+7 so substituting A=2B and C=B+7 into the 1st equation we get (2B)+B+(B+7)=57 so 4B=50 B=12.5 so A=2(12.5)=25 and C=12.5+7=19.5
So sides are 12.5,19.5,25 in length